OpenStudy (anonymous):

So when evaluating int 12cos^2(8x)dx I ended up with zero and that's not supposed to happen...can someone give me a hand?

OpenStudy (raden):

is this your question : 

OpenStudy (anonymous):

yes

OpenStudy (raden):

ok, first you can use the identity of trigono : cos^2 (a) = (1 + cos(2a))/2

OpenStudy (raden):

so, cos^2 (8x) = (1 + cos(2(8x)))/2 = (1 + cos(16x))/2, agree ?

OpenStudy (anonymous):

yeah I follow so far